# Supersymmetric BMS4 Algebras Revisited: Electric/Magnetic Superalgebras and Free Field Realization

### v1

@@ 1 Introduction
The study of asymptotic symmetries in general relativity originates from the seminal work of Bondi, van der Burg, Metzner, and Sachs [1, 2, 3, 4]. They demonstrated that the symmetry group of 44D asymptotically flat spacetime at null infinity extends beyond the Poincaré group. This enlarged symmetry is now referred to as the BMS4 algebra. A central feature of this algebra is the infinite-dimensional family of generators known as supertranslations, which extend the usual translations in Minkowski space. Beyond supertranslations, additional structure was subsequently identified. A systematic formulation of these “superrotations” was developed by Barnich and Troessaert [5, 6, 7, 8], who showed that they can be interpreted as local conformal transformations of the celestial sphere. In this sense, superrotations generalize global Lorentz transformations in the same way that supertranslations generalize translations. The resulting structure, often called the extended BMS algebra, takes the form
@@
While the BMS4 algebra encodes the asymptotic symmetries of flat spacetime, it is natural to ask whether these structures admit supersymmetric extensions. Supersymmetry plays a central role in modern theoretical physics, appearing in supergravity, string theory, and proposals for ultraviolet-complete quantum gravity theories. Extending BMS4 by incorporating fermionic generators enables the study of asymptotic symmetries in supergravity and the analysis of constraints imposed by supersymmetry at asymptotic infinities [37, 38, 39, 40]. Moreover, such extensions provide a systematic framework for examining celestial holography in supersymmetric regimes, offering insights into scattering amplitudes involving gravitinos and their superpartners [23, 41, 42, 43, 44]. The simplest extension introduces spin-$\frac{1}{2}$ generators $Q_{\alpha}$ alongside the bosonic supertranslation and superrotation generators [37, 45, 46, 47, 38, 48, 49, 50]. These generators satisfy anticommutation relations consistent with the Poincaré supersymmetry algebra, while commuting appropriately with the extended BMS generators. Schematically, the supersymmetric BMS algebra contains generators $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$222 $(l,\bar{l})$ for the super charge label the $\mathfrak{so}(3)$ spins in Lorentzian group $\mathfrak{so}(3)\times\mathfrak{so}(3)$. See Section 2.1 for more details. , and the commutation relation besides (1.1) is
@@
However, the above is not the full story of supersymmetric extension of the BMS4 algebra. In this work, we identify additional supersymmetric extensions, which can be classified as electric and magnetic types. Specifically, the electric super BMS4 algebra refers to the case where anticommutators of supercharges produce only supertranslations, while the magnetic super BMS4 algebra is characterized by anticommutators that also yield superrotations (See [51, 52] for early discussion on magnetic examples). In this classification, the simplest extension (1.2) corresponds to an electric super BMS4 algebra. This paper focuses on solutions admitting finite-dimensional subalgebras, which have potential applications in flat holography [53, 54, 23, 55]. For the electric case, there are ten distinct unextended superalgebras. It was realized in section 3 that only three of them can be realized without introducing spin higher than $\frac{1}{2}$ in the corresponding symmetry representations. The magnetic cases are more intricate. There are two types of chiral super BMS4 algebras: Type I (2.9ac) and Type II (2.9ad), along with four nonchiral superalgebras formed by combinations of these two types. Notably, for the nonchiral cases, $R$-symmetry is necessary for the closure of the algebra. The representation of the $R$-charges varies in each case, depending on the types of left- and right-handed supercharges. The appearance of the $R$-charges indicates that these combined super BMS4 algebras should be interpreted as extended supersymmetry.
@@
Beyond the algebraic formulation, this study presents an explicit free field realization of the super BMS4 algebra, by expressing the algebra generators as bilinears of the field modes. Specifically, we construct the supersymmetric extension of the massless electric scalar theory $\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi$, representing the simplest scalar multiplet. For the electric superalgebras, the theory (2.9y) admits three distinct electric superalgebras; conversely, for the magnetic superalgebras, each constructed theory is unique. Notably, although theory (2.9be) resembles the electric theory (2.9y), they are fundamentally different due to the distinct conformal dimensions of the fermionic fields. Interestingly, theory (2.9ca) establishes a relationship between the electric and magnetic scalars via $R$-symmetry, highlighting the essential role of magnetic scalars in supersymmetric flat holography. In theory (2.9dy), the emergence of a spin-$1$ massless boson suggests potential applications in supersymmetric soft theorems. Moreover, part of the theory (2.9dy) appeared in [21] as free perturbative supergravity theory. These realizations establish a direct correspondence between the algebraic structure of super BMS4 and physical observables, thereby facilitating the exploration of supersymmetric celestial conformal field theories [56] and the investigation of how supersymmetry influences the infrared behavior of gravitational scattering processes. Such theories serve as fundamental building blocks for understanding the spectrum and dynamics of supersymmetric theories in asymptotically flat spacetime. Moreover, the free field approach provides a systematic framework for deriving operator product expansions and correlation functions, which are essential for studying holographic dualities and soft theorems in supersymmetric gravity.
@@
The remainder of this paper is organized as follows. In Section 2, we discuss the possible super BMS4 algebras. After introducing the charge representations in Section 2.1, we identify ten electric super BMS4 algebras in Section 2.2, along with two chiral magnetic super BMS4 algebras and four nonchiral superalgebras in Section 2.3. We then proceed to the free field realizations of these algebras in Section 3. Specifically, in Section 3.2, we demonstrate that theory (2.9y) is trivial in the sense that it preserves three of the electric superalgebras. In the magnetic sector discussed in Section 3.3, we show that the field theories admitting different superalgebras are distinct. Lastly, we conclude this work in Section 4. Appendix A contains detailed calculations of the linear difference equations arising from the constraints on the representations.
@@ 2 Super BMS4 Algebras
It is well known that the global part of the BMS4 algebra is the 33D Carrollian conformal algebra [57]. As established in the literature, every aspect of Carrollian theories, such as the limiting procedures and types of field theories, has electric and magnetic versions [58, 57, 59]. In this paper, we adopt the terminology from Carrollian theory to classify the super BMS4 algebras. Specifically, we define a super BMS4 algebra as electric if the anticommutators of the supercharges yield only supertranslations. In contrast, we classify a super BMS4 algebra as magnetic if the anticommutators produce both superrotations and supertranslations:
@@ 2.1 Charge representation of BMS4 algebra
For these cases, $\{Q_{h-1},\dots,Q_{-h+1}\}$ forms a finite-dimensional representation space of the subalgebra $\{L_{1},L_{0},L_{-1}\}=\mathfrak{so}(3)$ with spin-l=h1l=h-1. In the remainder of this paper, we use the notation $\mathfrak{so}(3)$ spin-$l$ (and $\bar{l}$ for the anti-holomorphic part) to label the charge representations, extending this notation to general $l\in\mathbb{R}$.
@@
This is a linear difference equation which is only defined on discrete $n,r_{i},r_{j}\in\mathbb{Z}$ points. Numerically, we find the following useful solutions333 Here we consider solutions with $l\in\mathbb{Z}/2$ and $l>0$, i.e., charge representations that admit finite-dimensional sub-representations. See Appendix A for more details.
@@
$L_{m},M_{r,\bar{r}}]=\left(\frac{m}{2}-1*r\right)M_{m+r,\bar{r}}.$ (2.9k)
@@
$(l_{i})\otimes(l_{j})=(|l_{i}-l_{j}|)\oplus\cdots\oplus(l_{i}+l_{j}),$ (2.9l)
@@ 2.2 Electric super BMS4 algebra
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=[M_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{\bar{s}}]=0,\qquad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{\bar{r}}\}=M_{r,\bar{r}}.$ (2.9n)
@@
$\{L_{\pm 1,0},\bar{L}_{\pm 1,0},M_{\pm\frac{1}{2},\pm\frac{1}{2}},Q^{(\frac{1}{2},0)}_{\pm\frac{1}{2}},Q^{(0,\frac{1}{2})}_{\pm\frac{1}{2}}\}$, (2.9o)
@@
$J^{0,1}_{\text{P}}=\frac{i}{2}(-L_{1}-L_{-1}+\bar{L}_{1}+\bar{L}_{-1}),~J^{0,2}_{\text{P}}=\frac{1}{2}(L_{1}-L_{-1}+\bar{L}_{1}-\bar{L}_{-1}),~J^{0,3}_{\text{P}}=-L_{0}-\bar{L}_{0},$ (2.9p)
@@
However, the algebra obtained in (2.9n) does not exhaust all possible constructions of electric super BMS4. To proceed more systematically, we adopt a general strategy based on the following guiding principles. First, the supercharges are required to transform in charge representations, ensuring consistency with the underlying current algebra structure. Second, the labels $(l,\bar{l})$ of the supercharges are taken to lie in $\mathbb{Z}/2$ with $l,\bar{l}\geq 0$, so that the resulting algebra admits a finite-dimensional subalgebra, which can be viewed as a supersymmetric extension of the Poincaré algebra. Finally, to respect the spin–statistics theorem, the total spin must satisfy $l+\bar{l}\in\mathbb{Z}+\frac{1}{2}$. These ansatzes provide a systematic framework to classify and construct possible electric super BMS4 algebras beyond the minimal realization.
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{4},\frac{1}{4})}_{r,\bar{r}}\},\text{ with }\{Q^{(\frac{1}{4},\frac{1}{4})}_{r,\bar{r}},Q^{(\frac{1}{4},\frac{1}{4})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}},$ (2.9q)
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{4},\frac{3}{4})}_{r,\bar{r}}\},\text{ with }\{Q^{(\frac{3}{4},\frac{3}{4})}_{r,\bar{r}},Q^{(\frac{3}{4},\frac{3}{4})}_{s,\bar{s}}\}=\left(\frac{3}{4}r-\frac{3}{4}s\right)\left(\frac{3}{4}\bar{r}-\frac{3}{4}\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9r)
@@
We now turn to the case with two sets of supercharges. A representative example is provided by the extension of the super Poincaré algebra (2.9n), where both the holomorphic and anti-holomorphic labels satisfy the condition (2.9a). More generally, the supercharge $Q^{(\frac{1}{2},0)}_{r}$ can be extended to $Q^{(\frac{1}{2},0)}_{r,\bar{r}}$, with the original operator identified as $Q^{(\frac{1}{2},0)}_{r}=Q^{(\frac{1}{2},0)}_{r,0}$. This extension introduces additional super BMS4 charges while leaving the finite-dimensional subalgebra unchanged. In a completely analogous fashion, the supercharge $Q^{(0,\frac{1}{2})}_{\bar{r}}$ can be extended to $Q^{(0,\frac{1}{2})}_{r,\bar{r}}$. Following this logic, we obtain ten distinct realizations of the electric super BMS4 algebra, classified according to the rules (2.9), up to exchange of holomorphic and anti-holomorphic labels.
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}}.$ (2.9s)
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=\left(\frac{1}{2}\bar{r}-\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9t)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=\frac{3}{2}\bar{r}M_{r+s,\bar{r}+\bar{s}}.$ (2.9u)
holomorphic (2.9a), anti-holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.9v)

### v2

@@ 1 Introduction
The study of asymptotic symmetries in general relativity originates from the seminal work of Bondi, van der Burg, Metzner, and Sachs [1, 2, 3, 4]. They demonstrated that the symmetry group of 44D asymptotically flat spacetime at null infinity extends beyond the Poincaré group. This enlarged symmetry is now referred to as the BMS4 algebra. A central feature of this algebra is the infinite-dimensional family of generators known as supertranslations, which extend the usual translations in Minkowski space. Beyond supertranslations, additional structure was subsequently identified. A systematic formulation of these “superrotations” was developed by Barnich and Troessaert [5, 6, 7, 8], who showed that they can be interpreted as local conformal transformations of the celestial sphere. In this sense, superrotations generalize global Lorentz transformations in the same way that supertranslations generalize translations. The resulting structure, often called the extended BMS4 algebra, takes the form
@@
While the BMS4 algebra encodes the asymptotic symmetries of flat spacetime, it is natural to ask whether these structures admit supersymmetric extensions. Supersymmetry plays a central role in modern theoretical physics, appearing in supergravity, string theory, and proposals for ultraviolet-complete quantum gravity theories. Extending BMS4 by incorporating fermionic generators enables the study of asymptotic symmetries in supergravity and the analysis of constraints imposed by supersymmetry at asymptotic infinities [37, 38, 39, 40]. Moreover, such extensions provide a systematic framework for examining celestial holography in supersymmetric regimes, offering insights into scattering amplitudes involving gravitinos and their superpartners [23, 41, 42, 43, 44]. The simplest extension introduces spin-$\frac{1}{2}$ generators $Q_{\alpha}$ alongside the bosonic supertranslation and superrotation generators [37, 45, 46, 47, 38, 48, 49, 50]. These generators satisfy anticommutation relations consistent with the Poincaré supersymmetry algebra, while commuting appropriately with the extended BMS generators. Schematically, the supersymmetric BMS4 algebra contains generators $\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$222(l,l¯)($\bar{l}$) for the super charge label the $\mathfrak{so}(3)$ spins in Lorentzian group $\mathfrak{so}(3)\times\mathfrak{so}(3)$. See Section 2.1 for more details. , and the commutation relation besides (1.1) is
@@
In [51, 52, 53], the authors proposed the democratic (homogeneous) and despotic (inhomogeneous) limits to obtain two types of super BMS3 algebras. Applying the same method to the BMS4 case yields (1.2) as well as algebra (2.29) respectively. However, taking limit is not the full story of supersymmetric extension of the BMS4 algebra. In this work, we identify additional supersymmetric extensions, which can be classified as electric and magnetic types. Specifically, the electric super BMS4 algebra refers to the case where anticommutators of supercharges produce only supertranslations, while the magnetic super BMS4 algebra is characterized by anticommutators that also yield superrotations (See [54, 55] for early discussion on magnetic examples). In this classification, the simplest extension (1.2) corresponds to an electric super BMS4 algebra. This paper focuses on solutions admitting finite-dimensional subalgebras, which have potential applications in flat holography [56, 57, 23, 58]. For the electric case, there are ten distinct unextended superalgebras. It was realized in section 3 that only three of them can be realized without introducing spin higher than $\frac{1}{2}$ in the corresponding symmetry representations. The magnetic cases are more intricate. There are two types of chiral super BMS4 algebras: Type I (2.29) and Type II (2.30), along with four nonchiral superalgebras formed by combinations of these two types. Notably, for the nonchiral cases, $R$-symmetry is necessary for the closure of the algebra. The representation of the $R$-charges varies in each case, depending on the types of left- and right-handed supercharges. The appearance of the $R$-charges indicates that these combined super BMS4 algebras should be interpreted as extended supersymmetry.
@@
Beyond the algebraic formulation, this study presents an explicit free field realization of the super BMS4 algebra, by expressing the algebra generators as bilinears of the field modes. Specifically, we construct the supersymmetric extension of the massless electric scalar theory $\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi$, representing the simplest scalar multiplet. For the electric superalgebras, the theory (3.21) admits three distinct electric superalgebras; conversely, for the magnetic superalgebras, each constructed theory is unique. Notably, although theory (3.39) resembles the electric theory (3.21), they are fundamentally different due to the distinct conformal dimensions of the fermionic fields. Interestingly, theory (3.47) establishes a relationship between the electric and magnetic scalars via $R$-symmetry, highlighting the essential role of magnetic scalars in supersymmetric flat holography. In theory (3.69), the emergence of a spin-$1$ massless boson suggests potential applications in supersymmetric soft theorems. Moreover, part of the theory (3.69) appeared in [21] as free perturbative supergravity theory. These realizations establish a direct correspondence between the algebraic structure of super BMS4 and physical observables, thereby facilitating the exploration of supersymmetric celestial conformal field theories [59] and the investigation of how supersymmetry influences the infrared behavior of gravitational scattering processes. Such theories serve as fundamental building blocks for understanding the spectrum and dynamics of supersymmetric theories in asymptotically flat spacetime. Moreover, the free field approach provides a systematic framework for deriving operator product expansions and correlation functions, which are essential for studying holographic dualities and soft theorems in supersymmetric gravity.
@@
The remainder of this paper is organized as follows. In Section 2, we discuss the possible super BMS4 algebras. After introducing the charge representations in Section 2.1, we identify ten electric super BMS4 algebras in Section 2.2, along with two chiral magnetic super BMS4 algebras and four nonchiral superalgebras in Section 2.3. We then proceed to the free field realizations of these algebras in Section 3. Specifically, in Section 3.2, we demonstrate that theory (3.21) is trivial in the sense that it preserves three of the electric superalgebras. In the magnetic sector discussed in Section 3.3, we show that the field theories admitting different superalgebras are distinct. Lastly, we conclude this work in Section 4. Appendix A contains detailed calculations of the linear difference equations arising from the constraints on the representations.
@@ 2 Super BMS4 Algebras
It is well known that the global part of the BMS4 algebra is the 33D Carrollian conformal algebra [^3] In this paper, the $d$-dimensional Carrollian conformal algebra refers to the finite-dimensional algebra that is isomorphic to the $(d+1)$-dimensional Poincaré algebra, as well as to the global part of the BMSd algebra. [60]. As established in the literature, every aspect of Carrollian theories, such as the taking limit procedures and types of field theories, has electric and magnetic versions [61, 60, 62, 51, 52, 53]. In this paper, we adopt the terminology from Carrollian theory to classify the super BMS4 algebras. Specifically, we define a super BMS4 algebra as electric if the anticommutators of the supercharges yield only supertranslations. In contrast, we classify a super BMS4 algebra as magnetic if the anticommutators produce both superrotations and supertranslations:
@@ 2.1 Charge representation of BMS4 algebra
For these cases, $\{Q_{h-1},\dots,Q_{-h+1}\}$ forms a finite-dimensional representation space of the subalgebra $\{L_{1},L_{0},L_{-1}\}=\mathfrak{so}(3)$ with spin-l=h1l=h-1. In the remainder of this paper, we use the notation introduced in [63], and label the representations with $\mathfrak{so}(3)$ spin-$l$ (and $\bar{l}$ for the anti-holomorphic part), with $l(\bar{l})$ extended to general $\mathbb{R}$.
@@
This is a linear difference equation which is only defined on discrete $n,r_{i},r_{j}\in\mathbb{Z}$ points. Numerically, we find the following useful solutions [^4] Here we consider solutions with $l\in\mathbb{Z}/2$ and $l>0$, i.e., charge representations that admit finite-dimensional sub-representations. See Appendix A for more details.
@@
$L_{m},M_{r,\bar{r}}]=\left(\frac{m}{2}-1*r\right)M_{m+r,\bar{r}}.$ (2.11)
@@
$(l_{i})\otimes(l_{j})=(|l_{i}-l_{j}|)\oplus\cdots\oplus(l_{i}+l_{j}),$ (2.12)
@@ 2.2 Electric super BMS4 algebra
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=[M_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{\bar{s}}]=0,\qquad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{\bar{r}}\}=M_{r,\bar{r}}.$ (2.14)
@@
$\{L_{\pm 1,0},\bar{L}_{\pm 1,0},M_{\pm\frac{1}{2},\pm\frac{1}{2}},Q^{(\frac{1}{2},0)}_{\pm\frac{1}{2}},Q^{(0,\frac{1}{2})}_{\pm\frac{1}{2}}\}$, (2.15)
@@
$J^{0,1}_{\text{P}}=\frac{i}{2}(-L_{1}-L_{-1}+\bar{L}_{1}+\bar{L}_{-1}),~J^{0,2}_{\text{P}}=\frac{1}{2}(L_{1}-L_{-1}+\bar{L}_{1}-\bar{L}_{-1}),~J^{0,3}_{\text{P}}=-L_{0}-\bar{L}_{0},$ (2.16)
@@
However, the algebra obtained in (2.14) does not exhaust all possible constructions of electric super BMS4. To proceed more systematically, we adopt a general strategy based on the following guiding principles. First, the supercharges are required to transform in charge representations, ensuring consistency with the underlying current algebra structure. Second, the labels $(l,\bar{l})$ of the supercharges are taken to lie in $\mathbb{Z}/2$ with $l,\bar{l}\geq 0$, so that the resulting algebra admits a finite-dimensional subalgebra, which can be viewed as a supersymmetric extension of the Poincaré algebra. Finally, to respect the spin–statistics theorem, the total spin must satisfy $l+\bar{l}\in\mathbb{Z}+\frac{1}{2}$. These ansatzes provide a systematic framework to classify and construct possible electric super BMS4 algebras beyond the minimal realization.
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{4},\frac{1}{4})}_{r,\bar{r}}\},\text{ with }\{Q^{(\frac{1}{4},\frac{1}{4})}_{r,\bar{r}},Q^{(\frac{1}{4},\frac{1}{4})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}},$ (2.17)
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{4},\frac{3}{4})}_{r,\bar{r}}\},\text{ with }\{Q^{(\frac{3}{4},\frac{3}{4})}_{r,\bar{r}},Q^{(\frac{3}{4},\frac{3}{4})}_{s,\bar{s}}\}=\left(\frac{3}{4}r-\frac{3}{4}s\right)\left(\frac{3}{4}\bar{r}-\frac{3}{4}\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.18)
@@
We now turn to the case with two sets of supercharges. A representative example is provided by the extension of the super Poincaré algebra (2.14), where both the holomorphic and anti-holomorphic labels satisfy the condition (2.9a). More generally, the supercharge $Q^{(\frac{1}{2},0)}_{r}$ can be extended to $Q^{(\frac{1}{2},0)}_{r,\bar{r}}$, with the original operator identified as $Q^{(\frac{1}{2},0)}_{r}=Q^{(\frac{1}{2},0)}_{r,0}$. This extension introduces additional super BMS4 charges while leaving the finite-dimensional subalgebra unchanged. In a completely analogous fashion, the supercharge $Q^{(0,\frac{1}{2})}_{\bar{r}}$ can be extended to $Q^{(0,\frac{1}{2})}_{r,\bar{r}}$. Following this logic, we obtain ten distinct realizations of the electric super BMS4 algebra, classified according to the rules (2.9), up to exchange of holomorphic and anti-holomorphic labels.
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{r}+\bar{s}}.$ (2.19)
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=\left(\frac{1}{2}\bar{r}-\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.20)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=\frac{3}{2}\bar{r}M_{r+s,\bar{r}+\bar{s}}.$ (2.21)
holomorphic (2.9a), anti-holomorphic (2.9c):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.22)